Contact Corporation just paid a dividend of $1.50 per share. The company expects that the dividend will grow at a rate of 10% for the next two years. After year two it is expected that the dividend will decline at a rate of 3% indefinitely. If the required return is 12%, what is the value of a share of stock?
Answer #1

Last Dividend, D0 = $1.50

Growth rate for next 2 years is 10% and a constant growth rate (g) of 3% thereafter

D1 = $1.5000 * 1.10 = $1.6500
D2 = $1.6500 * 1.10 = $1.8150
D3 = $1.8150 * 1.03 = $1.86945

Required Return, rs = 12%

P2 = D3 / (rs - g)
P2 = $1.86945 / (0.12 - 0.03)
P2 = $1.86945 / 0.09
P2 = $20.7717

P0 = D1 / (1 + rs) + D2 / (1 + rs)^2 + P2 / (1 + rs)^2
P0 = $1.65/1.12 + $1.815/1.12^2 + $20.7717/1.12^2
P0 = $19.48

Therefore, value of the stock is $19.48
Answer #2

2.66 (price in the 2 years discounted)+ DIV T+1/ r-0.03/(r+1)^2 = 19.21

answered by: Ayyas